Your investment portfolio has 20,000 shares of Fairfax Paint, which has an expected return of 5.47 percent and a price of 6 dollars per share, and 30,000 shares of Litchfield Design, which has a price of 3.6 dollars per share. If your portfolio has an expected return of 9.73 percent, then what is the expected return for Litchfield Design? Answer as a rate in decimal format so that 12.34% would be entered as .1234 and 0.98% would be entered as .0098.
Answer #1

Total amount invested in the portfolio = (20,000*6) + (30,000*3.6) = $228,000

Weight of Investment in Fairfax Paint = $120,000 / $228,000 = 0.527

Weight of Investment in Litchfield Design= $108,000 / $228,000 = 0.473

Portfolio Return = (Weight of Investment in Fairfax Paint*Return on Fairfax Paint) + (Weight of Investment in Litchfield Design*Return on Litchfield Design)

9.73% = (0.527*5.47%) + (0.473*Return on Litchfield Design)

6.84731% = 0.473*Return on Litchfield Design

14.4763% =  Return on Litchfield Design